OpenStudy (anonymous):

Find an equation of the line containing the given pair of points (1/6,1/3) and (5/6,1)

OpenStudy (anonymous):

Solve the following for y and simplify:\[\frac{y-\frac{1}{3}}{x-\frac{1}{6}}=\frac{y-1}{x-\frac{5}{6}} \]\[y=x+\frac{1}{6} \]

OpenStudy (anonymous):

\[m=(y_2-y_1)∕(x_2- x_1 )\] m=(1-1/3)/(5/6-1/6)=(2/3)/(4/6)=2/3*6/4=1 m=1 \[y-y_1=m(x-x_1)\] y-1=1(x-5/6) y=x-5/6+1=x+1/6 y=x+1/6